OpenStudy (anonymous):

BIRD is a rhombus BI=4x+15 RD=3(x+17) IR=_____ Answers include:159, 36, 2, 23

OpenStudy (anonymous):

This is a rhombus... all sides are equal ... 4x + 15 = 3(x + 17) x = 36 4x + 15 = 159 <<< answer

OpenStudy (kainui):

So birds and pie are both quadrilaterals? I learn so much every day.